Please consider two investment alternatives. Each investment requires a $100,000 initial cash payment. The net. ome of each investment is as follows: Investment A: $110,000 each year for three years Investment B: $60,000 in year 1, $21 Factors to consider: inc Present Value of $1 at Compound Interes Year 6% 10% 12% 15% 20% 0943 0.909 0.893 0.870 0.833 2 0.890 0.826 0.797 0.756 0.694 3 0840 0.751 0.712 0.658 0.579 4 0.792 0.683 0.636 0.572 0.482 5 0.747 0.621 0.567 0.497 0.402 6 0.705 0.564 0.507 0.432 0.335 7 0.665 0.513 0.452 0.376 0.279 8 0.627 0.467 0.404 0.327 0.233 9 0.592 0.424 0.361 0.284 0.194 10 0.558 0.386 0.322 0.247 0.162 10,000 in year 2 and $60,000 in year 3 Depreciation is estimated at $10,000 per year The required rate of return is 10% 1. 2. 1. What is the Net Cash Flow from each investment? What is the Cash Payback Period from each investment? Which investment is preferred using the payback model? 2. 3. What is the Net Present Value of each investment? Which investment is preferred using NPV model?
Answer #1

1.Net Cash Flow:

Investment A , annual cash flow = $110,000+10,000 = $120,000 (since depreciation is a non cash expense)

Net Cash Flow = 120,000*PVAF(10%, 3years)

= 120,000*2.486

=$298,320

Investment B = 70,000*0.909 + 220,000*0.826 + 70,000*0.751

= $297,920

2. cash payback period:

Investment A: 100,000/120,000 = 0.833 years

B: 1 + 30,000/220,000 = 1+ 0.136 = 1.136 years

(70,000 in year 1, remaining 30,000 from year 2)

Preferred Investment = A, since payback period is lower for A

3. NPV = Present Value of Cash Inflows – Present Value of Cash Outflows

A: $298,320 - $100,000 = $198,320

B: $297,920 - $100,000 = $197,920

Preferred Investment = A., since higher NPV